OpenStudy (anonymous):

A monkey is given 12 blocks : 3 shaped like squares, 3 rectangles, 3 triagles and 3 circles. if it draws three of each kind in order-say 3 triangles,then 3 squares, and so on-would you suspect that the monkey associates identically shaped figures? Calculate the propability of this event.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

nice shoes!

OpenStudy (misty1212):

we can calculate the probability of 3 triangles, 3 squares, then 3 rectangles first then multiply that result by 6, since it could be TSR TRS STR SRT RTS RST

OpenStudy (misty1212):

first is a triangle \(\frac{3}{12}\) second is a triangle \(\frac{2}{11}\) third is a triangle \(\frac{1}{10}\) fourth is a rectangle \(\frac{3}{9}\) fifth is a rectangle \(\frac{2}{8}\) etc

OpenStudy (misty1212):

the pattern should be pretty clear when you multiply these you will get \[\frac{3\times 3!}{12!}\]

OpenStudy (misty1212):

since there are 6 ways to do it, as above, final answer (if i am reading it right) is \[\frac{6\times 3\times 3!}{12!}\]which is easy enough to compute

OpenStudy (misty1212):

can i get the shoes now?

OpenStudy (anonymous):

hahah thank you :) well i dont know about the shoes but you are sure getting best response :p

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

okay i am actually going through the problem and I dont think that is the solution :/